### v1

@@ 4.2 Star formation main sequence
One of the common indicators of the galaxy’s current star formation activity is its relation between SFR and stellar mass () commonly known as the star-forming main sequence (MS). The position of a galaxy compared to the MS helps in classifying it as either a star-burst galaxy (above the MS), a passive galaxy (below the MS), or as a normal SF galaxy (close to the MS; e.g. Elbaz et al. 2007; Noeske et al. 2007; Whitaker et al. 2012; Johnston et al. 2015; Davies et al. 2016; Siudek et al. 2018, Davies et al. submitted). The MS relation exists across a range of epochs and environments and is roughly linear, with the normalization increasing with redshift (e.g. Schreiber et al. 2015; Lee et al. 2015; Thorne et al. 2021) suggesting that the majority of star-forming galaxies are in a self-regulated equilibrium state (e.g. Bouché et al. 2010; Daddi et al. 2010; Genzel et al. 2010; Lagos et al. 2011; Lilly et al. 2013; Davé et al. 2013; Mitchell et al. 2016).
@@
In this paper, we demonstrate the MS relation for BGS galaxies. The BGS sample is a flux-limited survey at low redshift () divided into two programs: BRIGHT with and DARK reaching fainter galaxies at (Hahn et al., 2023b). We utilize the rest-framed colors derived with CIGALE: vs. to construct the UVJ diagram (Williams et al. 2009; Whitaker et al. 2012) used to identify red and blue galaxies. The UVJ diagram for a sample of 356,304 BGS (BRIGHT and DARK) galaxies999To select a sample of BGS galaxies we implement only the cuts outlined in Sec. 4.1, i.e. and . is shown in Fig. 4 with a separation cut following the definition given by Whitaker et al. (2012): , and 101010The cut is no longer implemented, as it is a false upper limit imposed on the quiescent population (van der Wel et al., 2014).. A gradual change of specific SFR (sSFR) with colors is clear. Red galaxies are characterized by low sSFR (), while blue galaxies are actively star forming ().
@@
The distribution of stellar mass and SFR for red and blue galaxies is shown in fig. 5. We find a similar number of red and blue galaxies (48%, and 52%, respectively) among the flux-limited BGS sample, with a mean stellar mass and 9.89 for red and blue galaxies, respectively (see left panel in Fig. 5). As expected, the massive red galaxies are characterized by lower SFRs than low-mass blue galaxies (with a mean SFR and 0.3 for red and blue galaxies, respectively; see right panel in Fig. 5). The blue galaxies tend to follow the MS according to Schreiber et al. (2015) with slightly higher SFR values possibly due to the choice of the dust attenuation prescription (see e.g. Siudek et al. 2018). The clear MS trend not only validates the SED fitting procedures in recreating the proper physical properties for a population of BGS galaxies but also showcases the utility of derived rest-framed colors for galaxy classification purposes. We note that to quantitatively compare the fraction and properties of red and blue galaxies one should take into account the selection biases. For example, selecting a mass complete sample (selected following prescription from Pozzetti et al. 2010) changes the fraction of red and blue galaxies to 68% and 32%, respectively.
@@
In Fig. 6 the MS trend of star-forming galaxies observed at a median redshift of the BGS sample (; the median redshift of the BGS sample) is reproduced following the prescription given by Schreiber et al. (2015). Starburst and passive galaxies are commonly selected as galaxies that deviate for more than dex, and dex from the MS, respectively (e.g. Elbaz et al. 2007; Rodighiero et al. 2011; Sargent et al. 2012; Whitaker et al. 2012; Buat et al. 2019; Donevski et al. 2020). The lower limits distinguishing passive galaxies correspond to the transition between red and blue galaxies selected based on the UVJ diagram. The bulk of blue BGS galaxies are located close to the MS, while the red BGS are found under the MS.
@@ 5 Comparison to other catalogs
In this Section, we compare our estimates of the physical properties of DESI EDR galaxies with the COSMOS catalog (Weaver et al., 2022), AGN-COSMOS catalogs (Suh et al. 2019 and Thorne et al. 2022b) and DEVILS catalog (Thorne et al., 2022b). In App. C we further compare our estimates with: i) SDSS and Extended Baryon Oscillation Spectroscopic Survey (eBOSS) Firefly VAC (SDSS(Firefly DR16); Comparat et al. 2017), ii) SDSS MPA-JHU DR8, after the Max Planck Institute for Astrophysics and Johns Hopkins University (SDSS(MPA-JHU); Kauffmann et al. 2003b; Brinchmann et al. 2004; Tremonti et al. 2004), iii) GALEX-SDSS-WISE Legacy Catalog X2111111https://salims.pages.iu.edu/gswlc/ (GSWLC; Salim et al. 2016, 2018), and iv) the DESI VAC of the stellar masses and emission lines presented by Zou et al. 2024. The summary of the comparison of stellar masses from different catalogs (including the ones described in App. C for simplicity) is presented in Table 4 using metrics described in Sec. 5.1 to quantify the differences. There exist several other DESI VACs with stellar mass estimates, such as i) FastSpecFit Spectral Synthesis and Emission-Line Catalog FastSpecFit 3.2 (Moustakas et al. 2023, Moustakas et al. in prep), and ii) The DESI PRObabilistic Value-Added Bright Galaxy Survey catalog (PROVABGS; Hahn et al. 2023a). Fastspecfit is a stellar continuum and emission-line fitting code optimized to model jointly DESI optical spectra and broadband photometry using physically motivated stellar continuum and emission-line templates. PROVABGS also models jointly DESI spectroscopy and photometry using state-of-the-art Bayesian approach and returns full posterior distributions of the galaxy properties. We report the metrics for these reference catalogs in Table 4, however, a more detailed comparison is a subject of future work. For comparison purposes, all the stellar mass estimates are recalculated, if necessary, to the cosmology and IMF used to create our VAC.
@@ 5.2 COSMOS Catalog
Here we compare our stellar mass estimates with the ones from the COSMOS2020 catalog (Weaver et al., 2022). The catalog includes sources down to observed over a 2 of the Cosmic Evolution Survey (COSMOS) field. The catalog comes in two independent versions: the CLASSIC, based on the traditional aperture photometry performed on the PSF-homogenized images, with the exception of IRAC images (Laigle et al., 2016), and the FARMER, which uses a new profile-fitting photometric extraction tool based on the tractor (Lang et al., 2016). The COSMOS2020 catalog provides a photo-z accuracy and outlier rate below 1% for bright galaxies (). The photo-z accuracy and outlier rate degrade to , and , respectively for the faintest galaxies (. The CLASSIC version includes 1.7 million galaxies with photometry from optical to NIR, while the FARMER version is limited to almost one million galaxies within the UltraVISTA footprint to provide izYJHKs images used to construct galaxy models. For both catalogs, stellar masses are derived with two independent tools, LePhare (Arnouts et al., 2002; Ilbert et al., 2006) and Eazy (Brammer et al., 2008). In this analysis, we compare the FARMER version limited to the LePhare stellar mass estimates (Weaver et al. 2022; the statistical comparison of the stellar masses from CLASSIC version is presented in Table 4). LePhare is a SED fitting code that uses a set of templates generated using Bruzual & Charlot (2003) models and assuming a Chabrier (2003) IMF. The SFH is described by an exponentially declining SFH and a delayed SFH assuming solar and half-solar metallicities. The dust attenuation is modeled with the Calzetti et al. 2000 law and a curve with a slope (see App. A of Arnouts et al. 2013) with color excess limited to 0.7. The AGN templates are not incorporated.
@@
We select 2,000 galaxies with high redshift accuracy (121212This cut ensures that the comparison between COSMOS2020’s photo-z-based physical parameters and DESI’s spec-z-based physical parameters is not dominated by the redshift estimation difference.) and with stellar mass error from COSMOS2020(FARMER). We find the negligible differences between our stellar masses and the ones from COSMOS2020(FARMER) given by the = 0.001 dex and NMAD = 0.176 (see Table 4 and Sec. 5.1 for definition of metrics). The is even smaller than the one found when comparing COSMOS(FARMER) with Zou et al. (2024) (). We note that such median differences are well below the median error of stellar mass estimates ( dex; see Table 4). Finding such a consistency between different codes, parametrization of the SED fitting codes, and different SED coverage confirms the robustness of the stellar mass estimates. The comparison of our stellar mass estimates with the ones from COSMOS2020(FARMER) catalog in three redshift bins is shown in Fig. 7. There is no clear dependence on redshift, however at , the clear bimodal distribution in stellar mass reveals smaller offset ( = -0.02 and NMAD = 0.18) for low-mass galaxies () than for high-mass galaxies (, = 0.12 and NMAD = 0.22). A Pearson correlation coefficient found for the comparison of stellar masses from our catalog with the ones from COSMOS2020(FARMER) (r = 0.96) indicates a very strong positive linear relationship between the stellar mass estimates.
@@
We also compared the SFR estimates (see Fig. 8) finding of -0.37 and NMAD of 0.62 with a median error of 0.31. The rest-framed magnitudes are also in good agreement, namely is 0.05 and NMAD = 0.18 for the rest-framed r magnitude.
@@ 5.2.1 COSMOS AGN
CIGALE returns several AGN properties such as AGN fraction (AGNFRAC), viewing angle (AGNPSY), and AGN luminosity (AGNLUM) defined as the total (disk, dust remitted and scattered) luminosity of the AGN. We compare these properties with a sample of 754 X-ray AGN drawn from the Chandra-COSMOS Legacy Survey (Suh et al., 2019).
@@
Suh et al. (2019) obtained the AGN luminosities, and stellar masses among others using the SED-fitting code (Calistro Rivera et al., 2016) to model near SEDs of X-ray selected COSMOS AGN. AGN SEDs were decomposed into a nuclear torus, a host galaxy, and a starburst component with an additional component of the big blue bump template in the UV-optical range for BL AGN. The host galaxy models are generated from Bruzual & Charlot (2003) SSP assuming solar metallicity, Chabrier (2003) IMF, and a simple exponentially declining SFH.
@@
Suh et al. (2019) provided stellar masses derived with which are dex lower than our estimates with CIGALE. A similar underestimation of stellar masses was found by Thorne et al. (2022b) when comparing stellar masses to DEVILS stellar mass estimates derived with PROSPECT (Davies et al., 2021; Thorne et al., 2021, 2022b, 2022a). Thorne et al. (2022b) discussed that this discrepancy is driven by the simplification of the host galaxy models (in the prescription of the SFH, metallicity, and dust) by the which focuses on recovering AGN properties. On the other hand, SED fitting codes such as PROSPECT and CIGALE are able to recreate the more complex nature of host galaxies.
@@
The AGN luminosities in our catalog stay in good agreement with that estimated by Suh et al. (2019) using , with a and NMAD of 0.378 despite the differences in the SED modeling. We also find close agreement with the bolometric AGN luminosity derived from the Chandra hard ( keV band) X-ray luminosities (see Fig. 9) characterized by a , NMAD of 0.648 and Pearson correlation of 0.85 independently on the redshift.
@@
Suh et al. (2019) distinguished the BL/unobscured and NL/obscured AGN types predetermined based on the optical properties (i.e based on the presence of the BL or NL in their spectra) or their photometric SED (i.e. whether is best fitted by an unobscured or obscured AGN template; see details in Marchesi et al. 2016 and Suh et al. 2019). We use this information to validate the AGNFRAC and AGNPSY derived with CIGALE. We find that 423 BL/unobscured AGN are characterized by higher AGN fraction (with a median AGNFRAC = 0.36) and larger viewing angle (AGNPSY = 67∘) than 397 NL/obscured AGN (with a median AGNFRAC = 0.27 and AGNPSY = 39∘). Only of BL/unobscured and NL/obscured AGN are observed with at least 2 MIR bands with SNR ¿ 3 suggesting that AGNFRAC and AGNPSY have the potential of discriminating between NL and BL AGN based on the CIGALE estimates even in the absence of the MIR information (see also Sec. 6.6). A more detailed discussion about AGN classification based on CIGALE is presented in Siudek et al. under DESI Collaboration review.
@@ 5.2.2 DEVILS Catalog
Thorne et al. (2022b) derived physical properties including stellar masses, SFR and AGN luminosities for DEVILS galaxies observed in the COSMOS field in using the PROSPECT SED fitting code incorporating AGN templates from Fritz et al. (2006) and a flexible star formation and metallicity (for details see Thorne et al. 2022b and also Thorne et al. 2021, 2022a). PROSPECT identified of BPT-selected AGN and derived AGN luminosities in close agreement with the luminosities derived from Chandra X-ray (Marchesi et al. 2016). The AGN identification based on the PROSPECT code is based on the AGN fraction, requiring AGNFRAC ¿ 0.1 (see also Table 2 in Thorne et al. 2022b for a comparison of PROSPECT AGN identification with standardly used techniques). We find PROSPECT counterparts for 5080 galaxies and we restrict our comparison to the sample of 1063 AGN and 2313 non-AGN galaxies with FIR photometry (FIRINPUT = 1131313Thorne et al. (2022b) found that for sources with lack of the FIR photometry AGN fraction is high (FRACAGN ¿ 0.8) and PROSPECT is not able to resolve properties of the host galaxy.) and AGN fraction AGNFRAC ¿ 0.1 and , respectively. We find the median difference on the stellar mass estimates to be on the level and NMAD for both AGN and non-AGN host galaxies. Restricting further the comparison to 386 AGN with WISE photometry (FLAGINFRARED ) we find the same AGN fraction ( = 0.27 and = 0.26), which further highlights the independence of the estimates on the code and prescription used for deriving properties of galaxies.
6 Physical properties: model & photometry dependence

### v2

@@ 4.2 Star formation main sequence
One of the common indicators of the galaxy’s current star formation activity is its relation between SFR and stellar mass (), commonly known as the star-forming main sequence (MS). The position of a galaxy compared to the MS helps in classifying it as either a star-burst galaxy (above the MS), a passive galaxy (below the MS), or as a normal SF galaxy (close to the MS; e.g., Elbaz et al. 2007; Noeske et al. 2007; Whitaker et al. 2012; Johnston et al. 2015; Davies et al. 2016; Siudek et al. 2018; Davies et al. 2025). The MS relation exists across a range of epochs and environments and is roughly linear, with the normalization increasing with redshift (e.g., Schreiber et al. 2015; Lee et al. 2015; Thorne et al. 2021) suggesting that the majority of star-forming galaxies are in a self-regulated equilibrium state (e.g., Bouché et al. 2010; Daddi et al. 2010; Genzel et al. 2010; Lagos et al. 2011; Lilly et al. 2013; Davé et al. 2013; Mitchell et al. 2016).
@@
In this paper, we demonstrate the MS relation for BGS galaxies. The BGS sample is a flux-limited survey at low redshift () divided into two programs: BRIGHT with and DARK reaching fainter galaxies at (Hahn et al., 2023b). We utilized the rest-frame colors derived with CIGALE: vs. to construct the UVJ diagram (Williams et al. 2009; Whitaker et al. 2012) used to identify red and blue galaxies. The UVJ diagram for a sample of 356,304 BGS (BRIGHT and DARK) galaxies111111To select a sample of BGS galaxies we implement only the cuts outlined in Sect. 4.1, i.e., and . is shown in Fig. 4 with a separation cut following the definition given by Whitaker et al. (2012): , and 121212The cut is no longer implemented, as it is a false upper limit imposed on the quiescent population (van der Wel et al., 2014).. A gradual change in the specific SFR (sSFR) with colors is evident. Red galaxies are characterized by low sSFR (), while blue galaxies are actively star forming ().
@@
The distribution of stellar mass and SFR for red and blue galaxies is shown in fig. 5. We find a similar number of red and blue galaxies (48%, and 52%, respectively) among the flux-limited BGS sample, with a mean stellar mass and 9.89 for red and blue galaxies, respectively (see left panel in Fig. 5). As expected, the massive red galaxies are characterized by lower SFRs than low-mass blue galaxies (with a mean SFR and 0.3 for red and blue galaxies, respectively; see right panel in Fig. 5). The blue galaxies tend to follow the MS according to Schreiber et al. (2015) with slightly higher SFR values possibly due to the choice of the dust attenuation prescription (see e.g., Siudek et al. 2018). The clear MS trend not only validates the SED fitting procedures in recreating the proper physical properties for a population of BGS galaxies but also showcases the utility of derived rest-frame colors for galaxy classification purposes. We note that to quantitatively compare the fraction and properties of red and blue galaxies one should take into account the selection biases. For example, selecting a mass complete sample (selected following prescription from Pozzetti et al. 2010) changes the fraction of red and blue galaxies to 68% and 32%, respectively.
@@
In Fig. 6, the MS trend of star-forming galaxies observed at a median redshift of the BGS sample (; the median redshift of the BGS sample) is reproduced following the prescription given by Schreiber et al. (2015). Starburst and passive galaxies are commonly selected as galaxies that deviate for more than dex, and dex from the MS, respectively (e.g., Elbaz et al. 2007; Rodighiero et al. 2011; Sargent et al. 2012; Whitaker et al. 2012; Buat et al. 2019; Donevski et al. 2020). The lower limits distinguishing passive galaxies correspond to the transition between red and blue galaxies selected based on the UVJ diagram. The bulk of blue BGS galaxies are located close to the MS, while the red BGS are found under the MS.
@@ 5 Comparison to other catalogs
In this section, we compare our estimates of the physical properties of DESI EDR galaxies with the COSMOS (Weaver et al., 2022), AGN-COSMOS (Suh et al. 2019 and Thorne et al. 2022b), and DEVILS (Thorne et al., 2022b) catalogs. In Appendix C, we further compare our estimates with: i) SDSS and Extended Baryon Oscillation Spectroscopic Survey (eBOSS) Firefly VAC (SDSS(Firefly DR16); Comparat et al. 2017), ii) SDSS MPA-JHU DR8, after the Max Planck Institute for Astrophysics and Johns Hopkins University (SDSS(MPA-JHU); Kauffmann et al. 2003b; Brinchmann et al. 2004; Tremonti et al. 2004), iii) GALEX-SDSS-WISE Legacy Catalog X2131313https://salims.pages.iu.edu/gswlc/ (GSWLC; Salim et al. 2016, 2018), and iv) DESI VAC of the stellar masses and emission lines presented by Zou et al. 2024. The summary of the comparison of stellar masses from different catalogs (including the ones described in Appendix C for simplicity) is presented in Table 4, using metrics described in Sect. 5.1 to quantify the differences. There exist several other DESI VACs with stellar mass estimates, such as i) FastSpecFit Spectral Synthesis and Emission-Line Catalog FastSpecFit 3.2 (Moustakas et al. 2023, Moustakas et al. in prep) and ii) The DESI PRObabilistic Value-Added Bright Galaxy Survey catalog (PROVABGS; Hahn et al. 2023a). Fastspecfit is a stellar continuum and emission-line fitting code optimized to jointly model DESI optical spectra and broadband photometry using physically motivated stellar continuum and emission-line templates. PROVABGS also models jointly DESI spectroscopy and photometry using state-of-the-art Bayesian approach and returns full posterior distributions of the galaxy properties. We report the metrics for these reference catalogs in Table 4, however, a more detailed comparison is the subject of a future work. For comparison purposes, all the stellar mass estimates are recalculated, if necessary, to the cosmology and IMF used to create our VAC.
@@ 5.2 COSMOS Catalog
Here, we compare our stellar mass estimates with the ones from the COSMOS2020 catalog (Weaver et al., 2022). The catalog includes sources down to observed over a 2 of the Cosmic Evolution Survey (COSMOS) field. The catalog comes in two independent versions: the CLASSIC, based on the traditional aperture photometry performed on the PSF-homogenized images, with the exception of IRAC images (Laigle et al., 2016), and FARMER, which uses a new profile-fitting photometric extraction tool based on the tractor (Lang et al., 2016). The COSMOS2020 catalog provides a photo-z accuracy and outlier rate below 1% for bright galaxies (). The photo-z accuracy and outlier rate degrade to , and , respectively, for the faintest galaxies (. The CLASSIC version includes 1.7 million galaxies with photometry from optical to NIR, while the FARMER version is limited to almost one million galaxies within the UltraVISTA footprint to provide izYJHKs images used to construct galaxy models. For both catalogs, stellar masses are derived with two independent tools: LePhare (Arnouts et al., 2002; Ilbert et al., 2006) and Eazy (Brammer et al., 2008). In this analysis, we compare the FARMER version limited to the LePhare stellar mass estimates (Weaver et al. 2022; the statistical comparison of the stellar masses from CLASSIC version is presented in Table 4). LePhare is a SED fitting code that uses a set of templates generated using Bruzual & Charlot (2003) models and assuming a Chabrier (2003) IMF. The SFH is described by an exponentially declining SFH and a delayed SFH assuming solar and half-solar metallicities. The dust attenuation is modeled with the Calzetti et al. 2000 law and a curve with a slope (see App. A of Arnouts et al. 2013) with color excess limited to 0.7. The AGN templates are not incorporated.
@@
We selected 2,000 galaxies with a high redshift accuracy (151515This cut ensures that the comparison between COSMOS2020’s photo-z-based physical parameters and DESI’s spec-z-based physical parameters is not dominated by the redshift estimation difference.) and a stellar mass error of from COSMOS2020(FARMER). We found the negligible differences between our stellar masses and the ones from COSMOS2020(FARMER) given by the = 0.001 dex and NMAD = 0.176 (see Table 4 and Sect. 5.1 for definitions for these metrics). The value is even smaller than the one found when comparing COSMOS(FARMER) with Zou et al. (2024) (). We note that such median differences are well below the median error of stellar mass estimates ( dex; see Table 4). Finding such a consistency between different codes, parametrization of the SED fitting codes, and different SED coverage confirms the robustness of the stellar mass estimates. The comparison of our stellar mass estimates with the ones from COSMOS2020(FARMER) catalog in three redshift bins is shown in Fig. 7. There is no clear dependence on redshift, however at , the clear bimodal distribution in stellar mass reveals smaller offset ( = -0.02 and NMAD = 0.18) for low-mass galaxies () than for high-mass galaxies (, = 0.12 and NMAD = 0.22). A Pearson correlation coefficient found for the comparison of stellar masses from our catalog with the ones from COSMOS2020(FARMER) (r = 0.96) indicates a very strong positive linear relationship between the stellar mass estimates.
@@
We also compared the SFR estimates (see Fig. 8) finding of -0.37 and NMAD of 0.62 with a median error of 0.31. The rest-frame magnitudes are also in good agreement, namely is 0.05 and NMAD = 0.18 for the rest-frame r magnitude.
@@ 5.2.1 COSMOS AGN
CIGALE returns several AGN properties such as the AGN fraction (AGNFRAC), viewing angle (AGNPSY), and AGN luminosity (AGNLUM) defined as the total (disk, dust remitted, and scattered) luminosity of the AGN. We compared these properties with a sample of 754 X-ray AGN drawn from the Chandra-COSMOS Legacy Survey (Suh et al., 2019).
@@
Suh et al. (2019) obtained the AGN luminosities and stellar masses (among other properties) using the SED-fitting code (Calistro Rivera et al., 2016) to model near SEDs of X-ray selected COSMOS AGNs. The AGN SEDs were decomposed into a nuclear torus, a host galaxy, and a starburst component with an additional component of the big blue bump template in the UV-optical range for BL AGN. The host galaxy models are generated from Bruzual & Charlot (2003) SSP assuming solar metallicity, Chabrier (2003) IMF, and a simple exponentially declining SFH.
@@
Suh et al. (2019) provided stellar masses derived with that are dex lower than our estimates with CIGALE. A similar underestimation of stellar masses was found by Thorne et al. (2022b) when comparing stellar masses to DEVILS stellar mass estimates derived with PROSPECT (Davies et al., 2021; Thorne et al., 2021, 2022b, 2022a). Thorne et al. (2022b) explained that this discrepancy is likely driven by the simplification of the host galaxy models (in the prescription of the SFH, metallicity, and dust) by the , which is focused on recovering AGN properties. On the other hand, SED fitting codes such as PROSPECT and CIGALE are able to recreate the more complex nature of host galaxies.
@@
The AGN luminosities in our catalog stay in good agreement with that estimated by Suh et al. (2019) using , with a and NMAD of 0.378 despite the differences in the SED modeling. We also find close agreement with the bolometric AGN luminosity derived from the Chandra hard ( keV band) X-ray luminosities (see Fig. 9) characterized by a , NMAD of 0.648 and Pearson correlation of 0.85 independently of the redshift.
@@
Suh et al. (2019) distinguished the BL/unobscured and NL/obscured AGN types predetermined based on the optical properties (i.e., based on the presence of the BL or NL in their spectra) or their photometric SED (i.e., whether is best fitted by an unobscured or obscured AGN template; see details in Marchesi et al. 2016 and Suh et al. 2019). We use this information to validate the AGNFRAC and AGNPSY derived with CIGALE. We find that 423 BL/unobscured AGN are characterized by higher AGN fraction (with a median AGNFRAC = 0.36) and larger viewing angle (AGNPSY = 67∘) than 397 NL/obscured AGN (with a median AGNFRAC = 0.27 and AGNPSY = 39∘). Only of BL/unobscured and NL/obscured AGN are observed with at least 2 MIR bands with S/N ¿ 3 suggesting that AGNFRAC and AGNPSY have the potential of discriminating between NL and BL AGN based on the CIGALE estimates even in the absence of the MIR information (see also Sect. 6.6). A more detailed discussion about AGN classification based on CIGALE is presented in Siudek et al. (under DESI Collaboration review).
@@ 5.2.2 DEVILS catalog
Thorne et al. (2022b) derived physical properties including stellar masses, SFR and AGN luminosities for DEVILS galaxies observed in the COSMOS field in the using the PROSPECT SED fitting code incorporating AGN templates from Fritz et al. (2006) and flexible star formation and metallicity parameters (for details, see Thorne et al. 2022b and also Thorne et al. 2021, 2022a). PROSPECT identified of BPT-selected AGNs and derived AGN luminosities in close agreement with the luminosities derived from Chandra X-ray (Marchesi et al. 2016). The AGN identification based on the PROSPECT code is based on the AGN fraction, requiring AGNFRAC ¿ 0.1 (see also Table 2 in Thorne et al. 2022b for a comparison of PROSPECT AGN identification with standardly used techniques). We found PROSPECT counterparts for 5080 galaxies and we restricted our comparison to the sample of 1063 AGN and 2313 non-AGN galaxies with FIR photometry (FIRINPUT = 1161616Thorne et al. (2022b) found that for sources lacking FIR photometry, the AGN fraction is high (FRACAGN ¿ 0.8) and PROSPECT is not able to resolve properties of the host galaxy.) and an AGN fraction of AGNFRAC ¿ 0.1 and , respectively. We find the median difference on the stellar mass estimates to be on the level and NMAD for both AGN and non-AGN host galaxies. Further restricting the comparison to 386 AGNs with WISE photometry (FLAGINFRARED ), we find the same AGN fraction ( = 0.27 and = 0.26). This further highlights the independence of the estimates on the code and prescription used for deriving properties of galaxies.
6 Physical properties: model and photometry dependence